OpenStudy (anonymous):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? Answer {-13, -3, 2, 3, 13} {12, 3, 0, 3, 13} {-17, -7, -2, 3, 13}

OpenStudy (anonymous):

for x=-3 y=5*-3-2=-15-2=-17 similarly plug the values of x from domain and find the range.

OpenStudy (anonymous):

Not understanding. I thought range was the smallest number from the largest number. Once that's determined (3-3)=0. Do I use the smallest to solve the equation and then the largest to solve? Lost and don't have my book.

OpenStudy (anonymous):

no domain is the values of xand range is the vlues obtained after pluugging the values of x e.g.,for x=-1 y=5*-1-2=-5-2=-7

OpenStudy (anonymous):

Going to try this right now and see what I get. Thanks so much!

OpenStudy (anonymous):

Got it! -17, -7, -2, 3, 13! Yeah! Just like to understand so that I can do more similar problems and I must show my work! Thanks bunches! It's actually easy!